OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 15 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column? PLEASE HELP FREE MEDALS !

OpenStudy (anonymous):

in 15 days you have half of 32 which is 16 in another 15 days you have half of 16, namely 8 in another 15 days you have half of 8, i.e. 4 in another 15 days... 2 and finally in 15 more days you have 1

OpenStudy (anonymous):

0 15 30 45 60 75 32 16 8 4 2 1

OpenStudy (anonymous):

it is a geometric sequence \[32,32\times \frac{1}{2},32\times \left(\frac{1}{2}\right)^2,...\]

OpenStudy (anonymous):

is that it for the ending ? is that how I answer @satellite73

OpenStudy (anonymous):

@CalebBeavers any ideas ?

OpenStudy (anonymous):

@satellite73 what will be my answer ? how will I express it ?

OpenStudy (whpalmer4):

Second column is a geometric sequence, as satellite73 demonstrated. What is the sequence in the first column? Isn't it just like counting? What sort of a sequence is counting?

OpenStudy (anonymous):

@whpalmer4 not sure / ? help

OpenStudy (anonymous):

finite?

OpenStudy (whpalmer4):

What kinds of sequences do you know about?

OpenStudy (anonymous):

Not many but im just confused as to that the first one is @whpalmer4

OpenStudy (whpalmer4):

Okay, if you don't know many, it shouldn't be hard to list them all, as I asked :-)

OpenStudy (whpalmer4):

How about arithmetic? Is counting like an arithmetic sequence? Is the first column of the table an arithmetic sequence? Aren't you just adding the same number over and over?